OpenStudy (anonymous):

how do I graph y^2+4x=0 using the (x-h)^2=4p(y-k)

OpenStudy (anonymous):

What do you mean exactly? Are you asking how you can put the first equation into the form you've seen for the second?

OpenStudy (anonymous):

yes i need to graph it and find the focus the directrix and p

OpenStudy (anonymous):

i would really appreciate the help I'm new haha

OpenStudy (anonymous):

Okay, just a sec...doing multiple things at once ;)

OpenStudy (anonymous):

I can be patient

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Your first and second equations are both parabolas, but they're different. The second equation will look like ones you're used to, whereas the first will open up differently. \[y^2+4x=0 \rightarrow y^2-4x\](i.e. subtracting 4x from both sides) puts your parabola in a form where you can read off things like focus and directrix. I'll continue in a sec.

OpenStudy (anonymous):

So, I made a mistake...great! It should be, \[y^2+4x=0 \rightarrow y^2=-4x\]

OpenStudy (anonymous):

Now, do me a favor and go to this website: www.wolframalpha.com and in the box at the beginning, type in: y^2+4x=0 Look at the graph and then come back. I'll continue with the focus and directrix.

OpenStudy (anonymous):

:) ty

OpenStudy (anonymous):

i think its suppose to be x= -y^2/4(1)

OpenStudy (anonymous):

because the focus is (1,0) and the directirx is (-1,0)

OpenStudy (anonymous):

but thank u for your help soo much it was lovely ;)

OpenStudy (anonymous):

The parabola you have is one that opens on its side. The form of this kind of parabola is\[(y-k)^2=4a(x-h)\]where a is your focus and (h,k) is the vertex. In your case, both h and k are zero, and if you compare the form with your equation, you'll see,\[y^2=4ax \rightarrow a=-1\] so your focus is at (-1,0) and directrix (1,0)

OpenStudy (anonymous):

hmm you have given me a lot to think about you have the opposite opinion of someone else i asked hmm ok ty :)

OpenStudy (anonymous):

You're welcome. Your equation above, x=-y^2/4(1) is the same as what I gave you...just looks different.

OpenStudy (anonymous):

thank you so much i really appreciate it :D